# Injective envelopes of $C^{*}$-algebras as maximal rigid multiplier covers

### v2

@@ Abstract.
canonically over $A$ by Hamana’s characterisation. In the commutative case $A=C(X)$, we recover Błaszczyk’s picture: for a maximal regular refinement $(X,T^{\ast})$ with irreducible identity, $G(X)=\beta(X,T^{\ast})$ and $I(C(X))\cong C(G(X))\cong M(C_{0}(U))$ for any dense open $U\subseteq G(X)$. Thus the paradigm maximise first, then compactify provides a unified conceptual framework on both sides of Gelfand duality.
@@
Institute of Mathematics, Czech Academy of Sciences (RVO: 67985840). Funding received from NCN Sonata-Bis 13 (2023/50/E/ST1/00067) is acknowledged with thanks.
@@ 1. Introduction
The injective envelope of a $C^{*}$-algebra, introduced by Hamana in his foundational papers [5, 6], has become a central tool in noncommutative analysis. For a compact Hausdorff $X$, injectivity on the commutative side amounts to extremal disconnectedness: $I(C(X))\cong C(K)$ with $K$ extremally disconnected. Gleason’s classical construction [4] of $K$ is intricate; Błaszczyk [1] later gave a strikingly concise route: first maximise the regular topology under an irreducibility constraint, then compactify to obtain $G(X)$. Indeed, in Błaszczyk argument, maximality is the driver and extremal disconnectedness is the consequence.
@@
Organisation. Section 2 recalls Hamana’s characterisation of $I(A)$ and the multiplier-ideal technology we need. Section 3 introduces multiplier covers, defines the forward order by AA‑preserving $\mathrm{ucp}$ maps, and proves that maximality forces essentiality via a lifting argument through $I(A)$. It also contains an intrinsic transfinite existence proof of maximal rigid covers using minimal $A^{1}$‑projections and the Choi-Effros product. The main identification $M(E_{\max})\cong I(A)$ is established in Section 4, and Section 5 revisits the commutative picture, recovering Błaszczyk. We end with remarks on functoriality and local multipliers.
@@ 2.1. Hamana’s injective envelope
Equivalently, $I(A)$ is the unique rigid essential extension of $A$: whenever $q:I(A)\to\mathsf{F}$ is a complete quotient map and $q|_{A}$ is a complete isometry, then $q$ is injective.
@@ 2.2. Multiplier algebras and quotienting ideals
We shall use the following standard fact.
@@
Let $E$ be a $C^{\ast}$-algebra and $I\lhd M(E)$ a closed two-sided ideal. Put
[…]
Then the quotient map $q_{K}:E\to E/K$ induces a surjective *-homomorphism
@@
whose kernel contains $I$. Conversely, for a closed ideal $K\lhd E$, the strict extension $M(E)\twoheadrightarrow M(E/K)$ has kernel $\{m\in M(E):mE\subseteq K,\ Em\subseteq K\}$.
@@
See [8, §3.12] and [2]. For $x\in I$ one has $xE,Ex\subseteq K$, hence $x\in\ker\pi_{K}$. The converse kernel description is the usual characterisation of multipliers vanishing modulo $K$. ∎
@@
If $B$ is unital and $J\lhd B$, the canonical map $B\to M(J)$ given by left/right multiplication is a surjective *-homomorphism with kernel $\mathrm{Ann}(J):=\{b\in B:bJ=Jb=0\}$; hence
$M(J)\ \cong\ B/\mathrm{Ann}(J).$
Thus $M(J)=B$ iff $J$ is essential in $B$. In particular, even if $B$ is a $W^{*}$-algebra (e.g. a product of factors), $M(J)$ need not be $W^{*}$, because $\mathrm{Ann}(J)$ need not be weak-closed. This is consistent with Hamana’s examples where $I(A)$ is monotone complete but not von Neumann [7].
@@ 3. Multiplier covers and maximality
Throughout this subsection we embed $A$ faithfully and nondegenerately into $\mathcal{B}(H)$ and write $A^{1}$ for its unitisation inside $\mathcal{B}(H)$. By an $A^{1}$-map we mean a unital completely positive map fixing $A^{1}$ pointwise. An $A^{1}$-projection is an idempotent $A^{1}$-map. We endow the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point-ultraweak topology. In what follows, the only continuity we use is:
$\Phi_{\alpha}\to\Phi\ \ \text{point-ultraweak}\quad\Longrightarrow\quad\Phi_{\alpha}\circ T\ \to\ \Phi\circ T\ \ \text{point-ultraweak}$
for each bounded linear $T:\mathcal{B}(H)\to\mathcal{B}(H)$; this is automatic because $(\Phi\mapsto\Phi\circ T)$ is continuous when evaluated on fixed vectors.
@@
Then there exists a ucp map $P_{\infty}:\mathcal{B}(H)\to\mathcal{B}(H)$ with $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$ such that, for every $\lambda_{0}\in\Lambda$,
@@
Equip the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point-ultraweak topology. By compactness of the unit ball of completely bounded maps, a subnet of $(P_{\lambda})_{\lambda}$ converges point-ultraweakly to a ucp map $P_{\infty}$; pointwise convergence preserves $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$. Fix $\lambda_{0}$ and $x\in\mathcal{B}(H)$. Using $P_{\lambda}\circ P_{\lambda_{0}}=P_{\lambda}$ for all $\lambda\geq\lambda_{0}$,
$P_{\infty}\big(P_{\lambda_{0}}x\big)\;=\;\lim_{\lambda}P_{\lambda}\big(P_{\lambda_{0}}x\big)\;=\;\lim_{\lambda}P_{\lambda}(x)\;=\;P_{\infty}(x),$
so $P_{\infty}\circ P_{\lambda_{0}}=P_{\infty}$. A symmetric limit using $P_{\lambda_{0}}\circ P_{\lambda}=P_{\lambda}$ gives $P_{\lambda_{0}}\circ P_{\infty}=P_{\infty}$. Hence $P_{\infty}$ is a lower bound. ∎
@@
$\Phi\preccurlyeq\Psi\quad\Longleftrightarrow\quad\Phi\circ\Psi=\Phi=\Psi\circ\Phi\quad(\text{equivalently, }\operatorname{Ran}\Phi\subseteq\operatorname{Ran}\Psi).$
Moreover, any such minimal $P$ is idempotent: $P^{2}=P$.
@@
Let $P$ be minimal and let $\Xi$ be any $A^{1}$-projection. Then $P\circ\Xi\circ P=P$.
@@
Put $F:=P\circ\Xi\circ P$. Then $F$ is an $A^{1}$-projection and $F\preccurlyeq P$, so by minimality $F=P$. ∎
@@
Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be the inclusion and set $\widetilde{Q}:=\kappa\circ Q\circ P$. Then $\widetilde{Q}$ is an $A^{1}$-projection, so by Lemma 3.4,
$P\ =\ P\circ\widetilde{Q}\circ P\ =\ P\circ\kappa\circ Q\circ P\ =\ Q\circ P,$
using $P\circ\kappa=\mathrm{id}_{R}$ and $P^{2}=P$. For $r=P(x)\in R$ we get $Q(r)=Q(P(x))=P(x)=r$. ∎
@@
With $P$ and $R=P(\mathcal{B}(H))$ as above, every ucp map $\theta:R\to R$ with $\theta|_{A^{1}}=\mathrm{id}$ equals $\mathrm{id}_{R}$.
@@
Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be inclusion and set $\Xi:=\kappa\circ\theta\circ P$. Then $\Xi$ is an $A^{1}$-projection and by Lemma 3.4,
$P\ =\ P\circ\Xi\circ P\ =\ P\circ\kappa\circ\theta\circ P\ =\ \theta\circ P.$
Thus for $r=P(x)\in R$ we have $\theta(r)=\theta(P(x))=P(x)=r$. ∎
@@
This is the Choi-Effros theorem [3, Thm. 3.1]. ∎
@@
Let $P$ be minimal and equip $E:=P(\mathcal{B}(H))$ with the Choi-Effros product. Then:
@@
(b) This is Proposition 3.6 applied to $R=E$.
@@
Construct $E:=P(\mathcal{B}(H))$ via a minimal $A^{1}$-projection and endow it with the Choi-Effros product as before. By Proposition 3.8, $A\subseteq E$ is rigid and essential, and $E$ is injective as an operator system. Let $J\triangleleft E$ be an essential ideal (take $J=E$ if $E$ is unital). Then $M(J)\cong E$ (Remark 2.3), and the inclusion $\iota:A\hookrightarrow M(J)$ yields a rigid $A$-multiplier cover $(J,\iota)$.
@@
so there exists an $A$-preserving $\mathrm{ucp}$ map $\Theta:M(J)\to M(F)$ with $\Theta\circ\iota=\jmath$. Since $M(J)\cong E$ is injective as an operator system, there is an $A$-preserving $\mathrm{ucp}$ map $\Sigma:M(F)\to M(J)$ with $\Sigma\circ\jmath=\iota$ (Arveson extension along $\jmath$ into $E$).
@@
hence $\Sigma\circ\Theta$ is a $\mathrm{ucp}$ endomorphism of $M(J)$ fixing $A$, so by rigidity of $A\subseteq M(J)$ we have $\Sigma\circ\Theta=\mathrm{id}_{M(J)}$. Similarly,
@@
Let $j:A\hookrightarrow I(A)$ be the Hamana embedding (Theorem 2.1). Since $\pi|_{A}$ is a complete isometry, $\pi(A)$ is completely order isomorphic to $A$ via $\pi|_{A}$. By injectivity of $I(A)$, there exists a ucp map $\Sigma:C\to I(A)$ extending $j\circ(\pi|_{A})^{-1}:\pi(A)\to I(A)$. Compose to obtain $\Theta:=\Sigma\circ\pi:B\to I(A)$, which is ucp and satisfies $\Theta\circ\iota=j$.
Choose any essential ideal $E^{\prime}\lhd I(A)$ (e.g. $E^{\prime}=I(A)$ if unital) and identify $M(E^{\prime})$ with $I(A)$. Then $(E^{\prime},j)$ is a rigid $A$-multiplier cover. The map $\Theta$ shows $(E,\iota)\preccurlyeq(E^{\prime},j)$. If $\pi$ were not injective, then $\Theta$ would not be injective, hence $(E^{\prime},j)$ would strictly dominate $(E,\iota)$, contradicting maximality. Thus $\pi$ is injective. ∎

### v3

@@ Abstract.
canonically over $A$ by Hamana’s characterisation. In the commutative case $A=C(X)$, we recover Błaszczyk’s picture: for a maximal regular refinement $(X,T^{\ast})$ with irreducible identity, $G(X)=\beta(X,T^{\ast})$ and $I(C(X))\cong C(G(X))\cong M(C_{0}(U))$ for any dense cozero set $U\subseteq G(X)$ (in particular, for any dense open $F_{\sigma}$ set, e.g. an increasing union of clopen sets). Thus the paradigm maximise first, then compactify provides a unified conceptual framework on both sides of Gelfand duality.
@@
IM CAS (RVO 67985840).
@@ 1. Introduction
The injective envelope of a $C^{*}$-algebra, introduced by Hamana in his foundational papers [5, 6], has become a central tool in noncommutative analysis. For a compact Hausdorff $X$, injectivity on the commutative side amounts to extremal disconnectedness: $I(C(X))\cong C(K)$ with $K$ extremally disconnected. Gleason’s classical construction [4] of $K$ is intricate; Błaszczyk [1] later gave a strikingly concise route: first maximise the regular topology under an irreducibility constraint, then compactify to obtain $G(X)$. Indeed, in Błaszczyk’s argument, maximality is the driver and extremal disconnectedness is the consequence.
@@
Organisation. Section 2 recalls Hamana’s characterisation of $I(A)$ and the multiplier–ideal technology we need (in the correct level of generality). Section 3 introduces multiplier covers, defines the forward order by AA‑preserving $\mathrm{ucp}$ maps, and proves that maximality forces essentiality via a lifting argument through $I(A)$. It also contains an intrinsic transfinite existence proof of maximal rigid covers using minimal $A^{1}$‑projections and the Choi–Effros product. The main identification $M(E_{\max})\cong I(A)$ is established in Section 4, and Section 5 revisits the commutative picture, recovering Błaszczyk. We end with remarks on functoriality and local multipliers.
@@ 2.1. Hamana’s injective envelope
Moreover, the Hamana embedding $j:A\hookrightarrow I(A)$ is essential in the following quotient sense: whenever $q:I(A)\to\mathsf{F}$ is a complete quotient map and $q|_{A}$ is a complete isometry, then $q$ is injective.
@@ 2.2. Multiplier algebras and quotienting ideals
We shall use the following standard fact in the form that is valid without additional hypotheses, and record the stronger surjectivity statement under a $\sigma$-unitality assumption.
@@
Let $E$ be a $C^{\ast}$-algebra and $K\lhd E$ a closed ideal. Then the quotient map $q_{K}:E\to E/K$ extends uniquely to a strictly continuous *-homomorphism
@@
whose kernel is $\{m\in M(E):mE\subseteq K,\ Em\subseteq K\}$. The range of $\pi_{K}$ is strictly dense in $M(E/K)$. If $E$ is $\sigma$-unital, then $\pi_{K}$ is surjective.
Moreover, if $I\lhd M(E)$ is a closed ideal and $K:=\overline{IE+EI}\lhd E$, then $I\subseteq\ker\pi_{K}$.
@@
See Pedersen [8, §3.12.8–3.12.10] and Busby [2]. The strict extension and kernel identification are standard; strict density of the range is general, and surjectivity holds when $E$ is $\sigma$-unital. The inclusion $I\subseteq\ker\pi_{K}$ follows from $xE,Ex\subseteq K$ for $x\in I$. ∎
@@
If $B$ is unital and $J\lhd B$, there is a canonical *-homomorphism
$\lambda_{B,J}:\ B\longrightarrow M(J),\qquad b\mapsto(L_{b},R_{b}),$
by left/right multiplication; its kernel is $\mathrm{Ann}(J):=\{b\in B:bJ=Jb=0\}$. The map $\lambda_{B,J}$ is injective iff $J$ is essential in $B$, but it is not surjective in general (indeed, very rarely; e.g. $J=\mathcal{K}(H)$ in its unitisation). In particular, even inside a $W^{*}$ ambient algebra there is no automatic von Neumann phenomenon at the level of multipliers.
@@ 3. Multiplier covers and maximality
Throughout this subsection we embed $A$ faithfully and nondegenerately into $\mathcal{B}(H)$ and write $A^{1}$ for its unitisation inside $\mathcal{B}(H)$. By an $A^{1}$-map we mean a unital completely positive map fixing $A^{1}$ pointwise. An $A^{1}$-projection is an idempotent $A^{1}$-map. We endow the space of maps $\mathcal{B}(H)\to\mathcal{B}(H)$ with the point–ultraweak topology. The only continuity we use is:
$\Phi_{\alpha}\to\Phi\ \ \text{point–ultraweak}\quad\Longrightarrow\quad\Phi_{\alpha}\circ T\ \to\ \Phi\circ T\ \ \text{point–ultraweak}$
for each bounded linear $T:\mathcal{B}(H)\to\mathcal{B}(H)$.
@@
Then there exists an $A^{1}$-projection $P_{\infty}:\mathcal{B}(H)\to\mathcal{B}(H)$ such that, for every $\lambda_{0}\in\Lambda$,
@@
By point–ultraweak compactness of the set of ucp maps $\mathcal{B}(H)\to\mathcal{B}(H)$ and closedness of the conditions “unital”, “completely positive”, and “fixes $A^{1}$ pointwise”, there exists a subnet $(P_{\lambda(\alpha)})_{\alpha}$ converging point–ultraweakly to a ucp map $\Phi$ with $\Phi|_{A^{1}}=\mathrm{id}_{A^{1}}$.
Fix $\lambda_{0}\in\Lambda$ and $x\in\mathcal{B}(H)$. For all sufficiently large $\alpha$ we have $\lambda(\alpha)\geq\lambda_{0}$, hence $P_{\lambda(\alpha)}\circ P_{\lambda_{0}}=P_{\lambda(\alpha)}=P_{\lambda_{0}}\circ P_{\lambda(\alpha)}$. Passing to the point–ultraweak limit yields
$\Phi\circ P_{\lambda_{0}}=\Phi=P_{\lambda_{0}}\circ\Phi.$
Now form the Cesàro averages $A_{N}:=\frac{1}{N}\sum_{k=0}^{N-1}\Phi^{k}$. By point–ultraweak compactness, choose a subnet $(A_{N(\beta)})_{\beta}$ converging point–ultraweakly to a ucp map $P_{\infty}$ with $P_{\infty}|_{A^{1}}=\mathrm{id}_{A^{1}}$.
Since $(I-\Phi)\circ A_{N}=\frac{1}{N}(I-\Phi^{N})$ and $\|\Phi^{N}\|\leq 1$, we have $(I-\Phi)\circ A_{N}\to 0$ point–ultraweakly, and similarly $A_{N}\circ(I-\Phi)\to 0$. Passing to the subnet limit gives
$\Phi\circ P_{\infty}=P_{\infty}=P_{\infty}\circ\Phi.$
Hence for every $k\geq 0$, $P_{\infty}\circ\Phi^{k}=P_{\infty}$ and $\Phi^{k}\circ P_{\infty}=P_{\infty}$.
To see idempotence, note that for each $N$,
$P_{\infty}\circ A_{N}=\frac{1}{N}\sum_{k=0}^{N-1}P_{\infty}\circ\Phi^{k}=P_{\infty}.$
Taking $N=N(\beta)$ and passing to the limit gives $P_{\infty}^{2}=P_{\infty}$, so $P_{\infty}$ is an $A^{1}$-projection.
Finally, from $\Phi\circ P_{\lambda_{0}}=\Phi=P_{\lambda_{0}}\circ\Phi$ we obtain by induction $\Phi^{k}\circ P_{\lambda_{0}}=\Phi^{k}=P_{\lambda_{0}}\circ\Phi^{k}$ for all $k$. Thus $A_{N}\circ P_{\lambda_{0}}=A_{N}=P_{\lambda_{0}}\circ A_{N}$ for all $N$, and passing to the subnet limit yields $P_{\infty}\circ P_{\lambda_{0}}=P_{\infty}=P_{\lambda_{0}}\circ P_{\infty}$. ∎
@@
$\Phi\preccurlyeq\Psi\quad\Longleftrightarrow\quad\Phi\circ\Psi=\Phi=\Psi\circ\Phi.$
@@
Let $P$ be an $A^{1}$-projection on $\mathcal{B}(H)$ and set $R:=P(\mathcal{B}(H))$. Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ denote the inclusion. If $Q:R\to R$ is a ucp idempotent with $Q|_{A^{1}}=\mathrm{id}$, then
$\widetilde{Q}\ :=\ \kappa\circ Q\circ P:\mathcal{B}(H)\to\mathcal{B}(H)$
is an $A^{1}$-projection and satisfies $\widetilde{Q}\preccurlyeq P$.
@@
The map $\widetilde{Q}$ is ucp and fixes $A^{1}$ pointwise since $P$ and $Q$ do and $\kappa$ is unital completely positive. Moreover, using $P\circ\kappa=\mathrm{id}_{R}$ and $Q^{2}=Q$,
$\widetilde{Q}^{2}=\kappa QP\,\kappa QP=\kappa Q(P\kappa)QP=\kappa QQP=\kappa QP=\widetilde{Q}$.
Finally,
$P\circ\widetilde{Q}=P\kappa QP=\kappa QP=\widetilde{Q},\qquad\widetilde{Q}\circ P=\kappa QP^{2}=\kappa QP=\widetilde{Q}$,
so $\widetilde{Q}\preccurlyeq P$. ∎
@@
Let $\kappa:R\hookrightarrow\mathcal{B}(H)$ be the inclusion and set $\widetilde{Q}:=\kappa\circ Q\circ P$. By Lemma 3.4, $\widetilde{Q}$ is an $A^{1}$-projection and $\widetilde{Q}\preccurlyeq P$. By minimality of $P$, we have $\widetilde{Q}=P$.
For $r\in R$ write $r=P(x)$ for some $x\in\mathcal{B}(H)$. Then
$Q(r)=Q(P(x))=P(x)=r,$
since $\widetilde{Q}(x)=\kappa(Q(P(x)))=P(x)$. Hence $Q=\mathrm{id}_{R}$. ∎
@@
With $P$ and $R=P(\mathcal{B}(H))$ as above, every $\mathrm{ucp}$ map $\theta:R\to R$ with $\theta|_{A^{1}}=\mathrm{id}$ equals $\mathrm{id}_{R}$.
@@
Consider the Cesàro averages $A_{N}:=\frac{1}{N}\sum_{k=0}^{N-1}\theta^{k}$ on $R$. By point-ultraweak compactness, a subnet $(A_{N(\beta)})_{\beta}$ converges point-ultraweakly to a ucp map $Q:R\to R$. Each $A_{N}$ fixes $A^{1}$ pointwise (since $\theta$ does), hence $Q|_{A^{1}}=\mathrm{id}$.
From $(I-\theta)\circ A_{N}=\frac{1}{N}(I-\theta^{N})$ and $\|\theta^{N}\|\leq 1$, we have $(I-\theta)\circ A_{N}\to 0$ point-ultraweakly, and similarly $A_{N}\circ(I-\theta)\to 0$. Passing to the subnet limit yields
$Q\circ\theta=Q=\theta\circ Q.$
In particular, $Q\circ\theta^{k}=Q$ for all $k\geq 0$, so for each $N$,
$Q\circ A_{N}=\frac{1}{N}\sum_{k=0}^{N-1}Q\circ\theta^{k}=Q.$
Taking $N=N(\beta)$ and passing to the limit gives $Q^{2}=Q$, so $Q$ is a ucp idempotent on $R$ fixing $A^{1}$. By Lemma 3.5, we get $Q=\mathrm{id}_{R}$.
Finally, $Q\circ\theta=Q$ with $Q=\mathrm{id}_{R}$ forces $\theta=\mathrm{id}_{R}$. ∎
@@
This is the Choi–Effros theorem [3, Thm. 3.1]. ∎
@@
Let $P$ be minimal and equip $E:=P(\mathcal{B}(H))$ with the Choi–Effros product. Then:
@@
(b) This follows from Proposition 3.6 applied to $R=E$.
@@
Construct $E:=P(\mathcal{B}(H))$ via a minimal $A^{1}$-projection and endow it with the Choi–Effros product as before. By Proposition 3.8, $A\subseteq E$ is rigid and essential, and $E$ is injective as an operator system. Since $P$ is unital, $E$ is unital, hence $M(E)=E$. Set
J:=E,J\ :=\ E,
and let $\iota:A\hookrightarrow E=M(E)=M(J)$ be the inclusion (it extends unital to $A^{1}\to E$), which is therefore nondegenerate. Thus $(J,\iota)$ is a rigid $A$-multiplier cover.
@@
so there exists an $A$-preserving $\mathrm{ucp}$ map $\Theta:M(J)=E\to M(F)$ with $\Theta\circ\iota=\jmath$. Since $E$ is injective as an operator system, there is an $A$-preserving $\mathrm{ucp}$ map $\Sigma:M(F)\to E$ with $\Sigma\circ\jmath=\iota$.
@@
hence $\Sigma\circ\Theta:E\to E$ is $\mathrm{ucp}$ and fixes $A$, so by rigidity of $A\subseteq E$ we have $\Sigma\circ\Theta=\mathrm{id}_{E}$. Similarly,
@@
Let $j:A\hookrightarrow I(A)$ be the Hamana embedding (Theorem 2.1). Since $\pi|_{A}$ is a complete isometry, the operator system $\pi(A)\subseteq C$ is completely order isomorphic to $A$ via $\pi|_{A}$. By injectivity of $I(A)$, there exists a ucp map $\Sigma:C\to I(A)$ extending $j\circ(\pi|_{A})^{-1}:\pi(A)\to I(A)$.
Set $\Theta:=\Sigma\circ\pi:B\to I(A)$. Then $\Theta$ is ucp and $\Theta|_{A}=j$, hence $(E,\iota)\preccurlyeq(I(A),j)$ (viewing $(I(A),j)$ as a rigid $A$-multiplier cover, since $I(A)$ is unital and rigid over $A$).
By maximality of $(E,\iota)$ among rigid covers, we must also have $(I(A),j)\preccurlyeq(E,\iota)$, so there exists an $A$-preserving ucp map $\Lambda:I(A)\to B$ with $\Lambda\circ j=\iota$. Then $(\Lambda\circ\Theta)|_{A}=\mathrm{id}_{A}$, so by rigidity of the cover $(E,\iota)$ we get $\Lambda\circ\Theta=\mathrm{id}_{B}$. In particular, $\Theta$ is injective.
Finally, $\ker\pi\subseteq\ker\Theta=\{0\}$, hence $\pi$ is injective. ∎